Session: Mar 2010 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Maccaferri at Thu Mar 18 21:30:48 2010) Success:

Onsala (Michael Lindqvist at Fri Mar 19 06:24:59 2010) Success:
Foggy. Otherwise no known problems.

Torun (K. Borkowski at Fri Mar 19 06:49:24 2010) Minor failures:
Tackling the problem of noise diode mentioned in the previous experiment feedback turned out to be very troublesome and timeconsuming for a few of our technicians (led by Genek). After a few hours of their hectic work the observations were started from 18:38 UTC and henceforth continued without known problems.

Westerbork (Geert Kuper at Fri Mar 19 09:44:28 2010) Success:
12 Telescopes in the array. No known problems.

Effelsberg (Uwe Bach at Fri Mar 19 11:19:59 2010) Minor failures:
Started 30 min too late, because of azimuth engine problems.

Yebes (SGE at Mon Mar 22 08:31:52 2010) Success:
Not known problems.

Noto (P. Cassaro at Fri Mar 26 09:34:43 2010) Success:

Jodrell_Bank (alastair gunn at Mon Mar 29 12:25:18 2010) Success:
No known problems.

Cambridge (alastair gunn at Mon Mar 29 12:27:09 2010) Success:
Kn on BBCs 1 and 2. No known problems.